I have a hard time with word problems and I can't seem to get the 2 equations right to do the elimination. can you help by doing the steps so I can understand how it is done.
A. The sales representative here tells you they also have two floor plans available, but they only have 38 homes still for sale. Write an equation that illustrates the situation. Use x and y to denote floor plan one and floor plan two repectively.
B. The representative tells you that floor plan one sells for $175,000, and floor plan two sells for $200,000. She also mentions that all the available houses combined are worth $$7,200,000. Write an equation that illustrates this situation. Use the same variables you used in Part a

Answer by ewatrrr(24785)


Hi,
A. 38 houses still available for sale with two floor plans: x(plan1) and y(plan2)
x + y = 38
B. three times as many homes available with the second floor plan than the first
substituting 3x for y
$175,000*x + $200,000*y = $7,200,000
Note: x = (38-y)
$175,000*(38-y) + $200,000*y = $7,200,000
25,000y = 550,000
y = 22 and x = 16 (38-22)